OpenStudy (anonymous):

Six times Jason's collection of books and one-third of Nathan's collection add up to 134 books. One-third of Jason's collection and Nathan's entire collection add up to 31 books.

Directrix (directrix):

Let J = number of Jason's books Let N = number of Jason's books See if you think these equations accurately describe the problem: 6*J + (1/3)*N = 134 (1/3) * J + N = 31

Directrix (directrix):

If so, solve those simultaneously.

OpenStudy (anonymous):

so i just divide those two number

Directrix (directrix):

You can solve them simultaneously by addition/subtraction or substitution.